OpenStudy (anonymous):

50 divided by 5million

OpenStudy (zzr0ck3r):

\(\frac{50}{5000000} = \frac{5}{500000} = \frac{1}{100000} = .00001\)

OpenStudy (anonymous):

yes that is correct :)